Find the derivative of the function. f(x) = √9+/A+ = F(x) = Need Help? Talk to a Tutor Watch It Read It
Answer #1

Here foo= 79+14+ = fore(9+(4+ety j3 Diff. 6 worit. x. flox)=(9+4+xyfot£{4+x+y+y)(0xy f(X) = {(9+(4+x+) { jż ( (4+x+; 3)(x3)